# A Bayesian Modeling Framework for Estimation and Ground Segmentation of Cluttered Staircases

### v1

@@ IV Staircase Estimation
In this section we describe our staircase estimation pipeline that utilizes our modeling framework in conjunction with an Extended Kalman Filter (EKF). The EKF fuses our belief, $bel(\boldsymbol{{}_{L}X}_{k})$ at time $k$ with the incoming staircase measurement probability, $p(\boldsymbol{{}_{L}Z}_{k+1})$ at time $k\!+\!1$, to update our belief $bel(\boldsymbol{{}_{L}X}_{k+1})$ as,
$bel(\boldsymbol{{}_{L}X}_{k+1})$ $=\eta\>p(\boldsymbol{{}_{L}Z}_{k+1})\!\!\int\displaylimits_{\boldsymbol{{}_{L}X}_{k}}\!p(\boldsymbol{{}_{L}\hat{X}}_{k+1|k})bel(\boldsymbol{{}_{L}X}_{k})\mathrm{d}{\boldsymbol{{}_{L}X}}$
Here, $\eta$ is a normalizing factor and $p(\boldsymbol{{}_{L}\hat{X}}_{k+1|k})$ is the probability of predicted staircase state. The overall pipeline is shown in Fig. 5. Further, we utilize the estimated staircase parameters to segment stair surfaces on a cluttered staircase.
@@ IV-A Stair-to-Stair Matching
After initialization, new measurements undergo data association to match them with the current staircase state. This employs the mahalanobis distance, which quantifies the similarity between a measurement and the existing state while considering uncertainty. Each stair measurement is classified as either matched, new preceding, or new succeeding. Matched stairs correspond directly to existing stairs in the state, while new preceding/succeeding stairs indicate those located below/above the current lowest/highest stair respectively.
Figure 5: Block diagram depicting the overall estimation pipeline
@@ IV-B Process Model - Predicting Staircase State
$\boldsymbol{\hat{\xi}}^{(i)}_{k+1|k}$ $=\frac{1}{3}\left({\boldsymbol{\bar{\xi}}}_{k}^{(i)}+\boldsymbol{\bar{\xi}}_{k}^{(i,i-1)}+\boldsymbol{\bar{\xi}}_{k}^{(i,i+1)}\right),$
$\mathbf{\hat{p}}^{(i)}_{s}$ $=\frac{1}{3}\left(\mathbf{\bar{p}}_{s}^{(i)}+\mathbf{\bar{p}}_{s}^{(i,i-1)}+\mathbf{\bar{p}}_{s}^{(i,i+1)}\right),$
$\mathbf{\hat{p}}^{(i)}_{e}$ $=\frac{1}{3}\left(\mathbf{\bar{p}}_{e}^{(i)}+\mathbf{\bar{p}}_{e}^{(i,i-1)}+\mathbf{\bar{p}}_{e}^{(i,i+1)}\right),$
@@
$\boldsymbol{\hat{\xi}}^{(i)}_{k+1|k}=\boldsymbol{\bar{\xi}}_{k}^{(i,u+1)},\;\;\;\;\;\mathbf{\hat{p}}^{(i)}_{s}=\mathbf{\bar{p}}_{s}^{(i,u+1)},\;\;\;\;\;\mathbf{\hat{p}}^{(i)}_{e}=\mathbf{\bar{p}}_{e}^{(i,u+1)},$
@@
$\boldsymbol{\hat{\xi}}^{(i)}_{k+1|k}=\boldsymbol{\bar{\xi}}_{k}^{(i,u+N)},\;\;\;\;\;\mathbf{\hat{p}}^{(i)}_{s}=\mathbf{\bar{p}}_{s}^{(i,u+N)},\;\;\;\;\;\mathbf{\hat{p}}^{(i)}_{e}=\mathbf{\bar{p}}_{e}^{(i,u+N)},$
@@
Covariance Prediction. The predicted covariance matrix ($\mathbf{{}_{L}\hat{\Sigma}_{k+1|k}}$) not only reflects the uncertainty propagated from the current state but also incorporates the uncertainty associated with the staircase parameters themselves. This is achieved through the $\mathbf{\mathcal{R}}$ matrix, which models the uncertainty in the staircase parameters $(\sigma_{h},\sigma_{d},\sigma_{w},\sigma_{\psi_{s}},\sigma_{\psi_{n}},\sigma_{\Delta\psi})$. Crucially $\mathbf{\mathcal{R}}$ can be user-selected, providing a mechanism to adjust the trust in the process model. It is computed as,
$\mathbf{{}_{L}\hat{\Sigma}_{k+1|k}}$ $=\mathbf{F_{X}}\;\mathbf{{}_{L}{\Sigma}_{k}}\;\mathbf{F^{T}_{X}}+\mathbf{F_{S}}\mathbf{\mathcal{R}}\mathbf{F^{T}_{S}},$ (25)
@@ IV-C Updating Staircase Estimate
Next, the endpoint state is updated through a maximization process that refines the endpoints of each stair to maximize its stair width. For each stair $i$, this process considers both the predicted endpoints from the process model ($\mathbf{\hat{p}}^{(i)}_{s},\mathbf{\hat{p}}^{(i)}_{e}$) and the measured endpoints ($\mathbf{\bar{p}}^{(i)}_{sm},\mathbf{\bar{p}}^{(i)}_{em}$), selecting the two endpoints that maximize the XY Euclidean distance. This ensures that if the new measurement reveals previously unseen portions of the stair, the endpoint state is updated accordingly. Finally, these refined endpoints are reprojected onto the updated infinite-line state ($\mathbf{{}_{L}X_{k+1}}$) to ensure consistency between the two representations. This reprojection step effectively refines the endpoint estimates by incorporating the updated stair line location obtained from EKF. The resulting staircase estimate integrates the refined infinite-line and endpoint states, providing an accurate and robust representation of the staircase structure.
(a) Cluttered Staircase Scene
(b) Generated crop-box for each stair overlaid on point cloud
(c) Points selected in the crop-box for plane fitting
(d) Final segmented stair surface points after plane-fitting
Figure 6: Stair surface segmentation process for a cluttered staircase
@@ IV-D Stair Surface Segmentation
For points within each crop-box, the algorithm further refines the segmentation by fitting a plane through the points using RANSAC. Crucially, this plane fitting incorporates a constraint ensuring that the identified plane remains parallel to the ground plane. This constraint leverages the inherent geometric property of stairs, where the tread surface is typically parallel to the ground. The inliers to this fitted plane are then classified as the stair tread surface, effectively separating them from any remaining clutter or outliers within the crop-box. This approach, illustrated in Fig. 6, enables robust segmentation of stair surfaces even in cluttered environments.
@@ V-A1 System Overview
Experiments were conducted both in simulation (using Isaac Sim) and real-world environments (using a Boston Dynamics Spot robot). The Spot robot was equipped with an NVIDIA Jetson AGX Orin, a Velodyne VLP-16 LiDAR for localization and mapping (LiPO [LiPo]). To enhance pointcloud density, points from the Spot’s onboard RealSense sensors were registered to the LiDAR-based SLAM map. In Isaac Sim, a Velodyne VLP-128 LiDAR mounted on a simulated Spot robot generated high-fidelity point cloud data. Data collection involved manually teleoperating Spot to navigate around the staircase, ensuring complete capture by the onboard sensors. Manual teleoperation was crucial because the challenging nature of cluttered staircases hindered Spot’s autonomous navigation.
@@ V-A2 Dataset
A comprehensive dataset of 23 distinct staircase examples was curated to evaluate the proposed system, encompassing a diverse range of configurations (size, length, curvature, clutter). Specifically, 10 examples featured staircases without any clutter, while the remaining 13 included various types of clutter, such as large boxes obstructing staircase visibility and smaller items on the steps, representing different levels of clutter. The diverse set of staircases includes those exceeding 2 meters in width, with more than 12 steps, curved staircases, and both ascending and descending configurations.
@@ V-A3 Baselines
To evaluate our approach, we compare it against two staircase estimation baselines adapted from prior work [sriganesh2023fast] and one segmentation baseline. The simple averaging (AVG) merging algorithm estimates the staircase state by averaging the current estimate of each stair’s endpoints with corresponding new measurements. The maximizing algorithm (MAX) is similar to the averaging algorithm, except it selects the endpoints that maximize the Euclidean distance between the start and end points, effectively maximizing the estimated stair width. For segmentation, we use a Cloth Simulation Filter [zhang2016csf] as a baseline. This method simulates a cloth draped over the estimated staircase region, conforming to the underlying geometry to approximate stair surfaces.
@@ V-B Results
Figure 7: Staircase estimation results for 2 different examples (a) Original staircase scenarios with curved staircase (top) and long cluttered staircase (bottom) (b) All input measurements (magenta) overlaid on the point cloud (c) Estimated staircase by using averaging algorithm [sriganesh2023fast] on the input measurements. (d) Estimated staircase by using maximizing algorithm on the input measurements (e) Estimated staircase by using our proposed framework on the input measurements (f) Points classified as stair surface using our segmentation algorithm
@@
AVG [sriganesh2023fast] MAX EKF (Ours)
Stair Height
RMSE (cm)
$1.5\pm 0.7$ $1.9\pm 1.0$ $\boldsymbol{0.4\pm 0.2}$
Stair Depth
RMSE (cm)
$2.4\pm 2.3$ $3.7\pm 1.8$ $\boldsymbol{1.5\pm 0.5}$
Stair Width
RMSE (cm)
$194.6\pm 234.9$ $30.0\pm 19.6$ $\boldsymbol{11.9\pm 16.4}$
Stair Curvature
Angle RMSE (deg)
$4.5\pm 13.1$ $2.7\pm 2.7$ $\boldsymbol{1.2\pm 0.7}$
Stair XY Location
RMSE (cm)
$4.8\pm 3.2$ $6.0\pm 4.9$ $\boldsymbol{3.3\pm 1.4}$
Stair Z Location
RMSE (cm)
$2.7\pm 1.7$ $2.8\pm 2.1$ $\boldsymbol{1.9\pm 1.2}$
Stair Orientation
RMSE (deg)
$1.5\pm 0.86$ $2.2\pm 1.78$ $\boldsymbol{1.2\pm 0.7}$
TABLE I: Comparison of staircase estimation RMSE
As shown in TABLE I, our proposed method demonstrates significantly improved accuracy in estimating both the parameters and location of staircases with reductions in RMSE across all metrics. Notably, we achieve a 94% reduction in width error and 73% reduction in height error compared to the averaging approach [sriganesh2023fast]. Furthermore, our method achieves approximately a 30% reduction in RMSE for stair location (XY and Z). While the maximizing algorithm addresses some limitations of this prior work our approach consistently yields the most accurate results. This superior performance is attributed to our method’s ability to effectively incorporate information from the entire staircase structure, which helps to handle noise and incomplete measurements.
TABLE II presents the accuracy, precision, and recall of our proposed method for stair surface segmentation, compared against the Cloth Simulation Filter (CSF) [zhang2016csf] algorithm. Our method consistently outperforms CSF across all metrics. This improvement stems from our method’s ability to leverage information about the staircase structure, enabling more accurate differentiation between stair surface and flat debris. In contrast, CSF is purely geometric and often misclassifies such debris as ground. While CSF can be tuned to improve performance, we observed inconsistent results across different cluttered staircases with the same parameters. Our approach consistently provided robust and accurate segmentation regardless of the scenario.
Accuracy (%) Precision (%) Recall (%)
@@
Figure 9: Plot of execution times vs number of steps for the stair surface segmentation algorithm
To verify real-time performance, we evaluated our C++ implementation (using Eigen) on an NVIDIA Jetson AGX Orin. We analyzed the computational performance of both our EKF pipeline, model prediction and matching algorithm, as well as the stair surface segmentation. Fig. 8 illustrates the performance of the EKF (shown in red) and the model prediction/matching (shown in blue) components, plotting the total computation time as a function of the number of stairs in the staircase (ranging from 3 to 20). Both the prediction and matching times increase with the number of stairs, with the EKF component exhibiting a steeper rise. Fig. 9 illustrates the time required for stair surface segmentation, with times normalized for a point cloud of 250k points to account for varying point cloud densities across different staircases. Similar to the EKF, the computation time for segmentation shows an upward trend as the number of steps increases, reflecting the increased time required for a larger number of points. Despite these trends, the absolute segmentation times remain manageable, even for staircases with 20 steps. Crucially, the total time taken by both the estimation and segmentation methods, even for large staircases, remains under 30ms. This is well within 50ms budget to achieve our real-time execution target rate of 20Hz, confirming the feasibility of real-time operation.

### v2

@@ IV Staircase Estimation
Figure 4: Block diagram depicting the overall estimation pipeline.
In this section we describe our staircase estimation pipeline that utilizes our modeling framework. Our pipeline operates within a Bayesian framework: we begin with a belief about the staircase state at time $k$, $bel(\boldsymbol{{}_{L}X}_{k})$. This belief is then updated at time $k+1$ to $bel(\boldsymbol{{}_{L}X}_{k+1})$ by fusing it with the incoming staircase measurement likelihood, $p(\boldsymbol{{}_{L}Z}_{k+1})$, using the Bayesian update equation:
$bel(\boldsymbol{{}_{L}X}_{k+1})$ $=\beta\>p(\boldsymbol{{}_{L}Z}_{k+1})\!\!\int\displaylimits_{\boldsymbol{{}_{L}X}_{k}}\!p(\boldsymbol{{}_{L}\hat{X}}_{k+1|k})bel(\boldsymbol{{}_{L}X}_{k})\mathrm{d}{\boldsymbol{{}_{L}X}}$
Here, $\beta$ is a normalizing factor and $p(\boldsymbol{{}_{L}\hat{X}}_{k+1|k})$ represents the state transition probability obtained by the process model. Assuming Gaussian noise throughout our system, this Bayesian update could ideally be computed using a standard Kalman Filter [19]. However, as our process model is non-linear, we employ the Extended Kalman Filter (EKF) with a linear approximation of the process model. The overall pipeline is shown in Fig. 4. Further, we utilize the estimated staircase parameters to segment stair surfaces on a cluttered staircase.
@@ IV-A Stair-to-Stair Matching
After initialization, new measurements undergo data association to match them with the current staircase state. This employs the Mahalanobis distance, which quantifies the similarity between a measurement and the existing state while considering uncertainty. Each stair measurement is classified as either matched, new preceding, or new succeeding. Matched stairs correspond directly to existing stairs (within $3\sigma$) in the state, while new preceding/succeeding stairs indicate those located below/above the current lowest/highest stair respectively.
@@ IV-B Process Model - Predicting Staircase State
$\boldsymbol{\hat{\xi}}^{(i)}_{k+1|k}$ $=\frac{1}{3}\left(\boldsymbol{\xi}_{k}^{(i)}+\boldsymbol{\xi}_{k}^{(i,i-1)}+\boldsymbol{\xi}_{k}^{(i,i+1)}\right),$
$\mathbf{\hat{p}}^{(i)}_{s}$ $=\frac{1}{3}\left(\mathbf{p}_{s}^{(i)}+\mathbf{p}_{s}^{(i,i-1)}+\mathbf{p}_{s}^{(i,i+1)}\right),$
$\mathbf{\hat{p}}^{(i)}_{e}$ $=\frac{1}{3}\left(\mathbf{p}_{e}^{(i)}+\mathbf{p}_{e}^{(i,i-1)}+\mathbf{p}_{e}^{(i,i+1)}\right),$
@@
$\boldsymbol{\hat{\xi}}^{(i)}_{k+1|k}=\boldsymbol{\xi}_{k}^{(i,u+1)},\;\;\;\;\;\mathbf{\hat{p}}^{(i)}_{s}=\mathbf{p}_{s}^{(i,u+1)},\;\;\;\;\;\mathbf{\hat{p}}^{(i)}_{e}=\mathbf{p}_{e}^{(i,u+1)},$
@@
$\boldsymbol{\hat{\xi}}^{(i)}_{k+1|k}=\boldsymbol{\xi}_{k}^{(i,u+N)},\;\;\;\;\;\mathbf{\hat{p}}^{(i)}_{s}=\mathbf{p}_{s}^{(i,u+N)},\;\;\;\;\;\mathbf{\hat{p}}^{(i)}_{e}=\mathbf{p}_{e}^{(i,u+N)},$
@@
Covariance Prediction. The predicted covariance matrix ($\mathbf{{}_{L}\hat{\Sigma}_{k+1|k}}$) not only reflects the uncertainty propagated from the current state but also incorporates the uncertainty associated with the staircase parameters themselves. This is achieved through the $\boldsymbol{\mathcal{R}}$ matrix, which models the uncertainty in the staircase parameters $(\sigma_{h},\sigma_{d},\sigma_{w},\sigma_{\psi_{s}},\sigma_{\psi_{n}},\sigma_{\Delta\psi})$. Crucially $\boldsymbol{\mathcal{R}}$ can be user-selected, providing a mechanism to adjust the trust in the process model. Higher values in $\boldsymbol{\mathcal{R}}$ indicate less trust, causing the filter to rely more on incoming measurements.
[…]
@@ IV-C Updating Staircase Estimate
Next, the endpoint state is updated through a maximization process that refines the endpoints of each stair to maximize its stair width. For each stair $i$, this process considers both the predicted endpoints from the process model ($\mathbf{\hat{p}}^{(i)}_{s},\mathbf{\hat{p}}^{(i)}_{e}$) and the measured endpoints ($\mathbf{p}^{(i)}_{sm},\mathbf{p}^{(i)}_{em}$), selecting the two endpoints that maximize the XY Euclidean distance. This ensures that the endpoint state is updated upon observing previously unseen stair segments. Finally, these refined endpoints are re-projected onto the updated infinite-line state ($\mathbf{{}_{L}X_{k+1}}$) to ensure consistency between the two representations. This re-projection effectively refines the endpoint estimates by incorporating the updated stair line location obtained from EKF. The resulting staircase estimate integrates the refined infinite-line and endpoint states, providing an accurate and robust representation of the staircase structure.
(a) Cluttered Staircase Scene.
(b) Generated crop-box for each stair overlaid on point cloud.
(c) Points selected in the crop-box for plane fitting.
(d) Final segmented stair surface points after plane-fitting.
Figure 5: Stair surface segmentation process for a cluttered staircase.
@@ IV-D Stair Surface Segmentation
For points within each crop-box, the algorithm further refines the segmentation by fitting a plane through the points using RANSAC. Crucially, this plane fitting incorporates a constraint ensuring that the identified plane remains parallel to the ground plane. This constraint leverages the inherent geometric property of stairs, where the tread surface is typically parallel to the ground. The inliers to this fitted plane are then classified as the stair tread surface, effectively separating them from any remaining clutter or outliers within the crop-box. This approach, illustrated in Fig. 5, enables robust segmentation of stair surfaces even in cluttered environments. As our staircase state excludes the initial ground plane, it is not segmented in our algorithm. Incorporating this ground plane is left as future work.
@@ V-A1 System Overview
Experiments were conducted both in simulation (using Isaac Sim) and real-world environments (using a Boston Dynamics Spot robot). The Spot robot was equipped with an NVIDIA Jetson AGX Orin, a Velodyne VLP-16 LiDAR for localization and mapping (LiPO [20]). To enhance pointcloud density, points from the Spot’s onboard RealSense sensors were registered to the LiDAR-based SLAM map. In Isaac Sim, a Velodyne VLP-128 LiDAR mounted on a simulated Spot robot generated high-fidelity point cloud data. Data collection involved manually teleoperating Spot to navigate around the staircase, ensuring complete capture by the onboard sensors. Manual teleoperation was crucial because the challenging nature of cluttered staircases hindered Spot’s autonomous navigation.
@@ V-A2 Dataset
A comprehensive dataset of 23 distinct staircase examples (17 real-world, 6 simulation) was curated to evaluate the proposed system, encompassing a diverse range of configurations (size, length, curvature, clutter). Specifically, 10 examples featured staircases without any clutter, while the remaining 13 included various types of clutter, such as large boxes obstructing staircase visibility and smaller items on the steps, representing different levels of clutter. It included staircases with widths ranging between 1 and 10 meters and step counts ranging from 4 to 20. These included both ascending and descending orientations as well as curved and open-rise staircases.
@@ V-A3 Baselines
To evaluate our approach, we compare it against two staircase estimation baselines adapted from prior state-of-the-art [8] and one segmentation baseline. The simple averaging (AVG) merging algorithm estimates the staircase state by averaging the current estimate of each stair’s endpoints with corresponding new measurements. The maximizing algorithm (MAX) is similar to the averaging algorithm, except it selects the endpoints that maximize the Euclidean distance between the start and end points, effectively maximizing the estimated stair width. For segmentation, we use a Cloth Simulation Filter [15] as a baseline. This method simulates a cloth draped over the estimated staircase region, conforming to the underlying geometry to approximate stair surfaces.
@@ V-B Results
Figure 6: Staircase estimation results for 2 different examples (a) Original staircase scenarios with curved staircase (top) and long cluttered staircase (bottom) (b) All input measurements (magenta) overlaid on the point cloud (c) Estimated staircase by using averaging algorithm [8] on the input measurements. (d) Estimated staircase by using maximizing algorithm on the input measurements (e) Estimated staircase by using our proposed framework on the input measurements (f) Points classified as stair surface using our segmentation algorithm
R/S AVG [8] MAX EKF (Ours)
Stair Height RMSE (cm) R $1.8\pm 0.8$ $2.3\pm 0.9$ $\boldsymbol{0.6\pm 0.3}$
S $1.1\pm 0.5$ $1.0\pm 0.7$ $\boldsymbol{0.3\pm 0.1}$
Stair Depth RMSE (cm) R $2.0\pm 1.3$ $3.8\pm 2.1$ $\boldsymbol{1.3\pm 0.8}$
S $3.6\pm 4.0$ $3.5\pm 0.9$ $\boldsymbol{2.0\pm 1.5}$
Stair Width RMSE (cm) R $113.7\pm 112.5$ $29.9\pm 18.8$ $\boldsymbol{12.0\pm 18.7}$
S $424.0\pm 341.8$ $30.7\pm 23.7$ $\boldsymbol{11.8\pm 8.8}$
Stair Curvature RMSE (cm) R $2.0\pm 1.1$ $3.3\pm 2.8$ $\boldsymbol{1.5\pm 0.8}$
S $11.5\pm 26.1$ $1.4\pm 1.7$ $\boldsymbol{0.8\pm 0.7}$
Stair Location XY RMSE (cm) R $5.0\pm 3.4$ $6.8\pm 5.5$ $\boldsymbol{3.6\pm 1.5}$
S $4.0\pm 2.5$ $3.8\pm 1.8$ $\boldsymbol{2.9\pm 0.9}$
Stair Location Z RMSE (cm) R $3.3\pm 1.7$ $3.4\pm 2.1$ $\boldsymbol{2.3\pm 1.2}$
S $1.4\pm 0.7$ $1.2\pm 0.6$ $\boldsymbol{1.0\pm 0.9}$
Stair Orientation RMSE (deg) R $1.7\pm 0.8$ $2.6\pm 1.9$ $\boldsymbol{1.5\pm 0.7}$
S $0.9\pm 1.0$ $1.0\pm 1.1$ $\boldsymbol{0.7\pm 0.6}$
TABLE I: Comparison of staircase estimation RMSE. R indicates real-world scenarios, and S indicates simulation.
@@
As shown in TABLE I, our proposed method demonstrates significantly improved accuracy in estimating both the parameters and location of staircases across all metrics in simulation and in real-world. For real-world staircases, we achieve a 89% reduction in width error and 67% reduction in height error compared to the averaging approach [8]. Furthermore, our method achieves approximately a 30% reduction in RMSE for stair location (XY and Z). While the maximizing algorithm addresses some limitations of this prior work, our approach consistently yields the most accurate results. It’s worth noting that in simulation, the width error for the averaging method is considerably higher due to the presence of exceptionally wide ($10m$) staircases. Our superior performance is attributed to our method’s ability to effectively incorporate information from the entire staircase structure, which helps to handle noise and incomplete measurements.
Metric
@@
TABLE II presents the accuracy, precision, and recall of our proposed method for stair surface segmentation, compared against the Cloth Simulation Filter (CSF) [15]. Our method consistently outperforms CSF across all metrics. This improvement stems from our method’s ability to leverage information about the staircase structure, enabling more accurate differentiation between stair surface and flat debris. In contrast, CSF is purely geometric and often misclassifies such debris as ground. While CSF can be tuned to improve performance, we observed inconsistent results across different cluttered staircases with the same parameters. Our approach consistently provided accurate segmentation across all scenarios.
@@
Figure 8: Plot of execution times vs number of steps for the stair surface segmentation algorithm. Error bars indicate standard deviation.
To verify real-time performance, we evaluated our C++ implementation on an NVIDIA Jetson AGX Orin. We analyzed the computational performance of both our EKF pipeline, model prediction and matching algorithm, as well as the stair surface segmentation. Fig. 7 illustrates the performance of the EKF (shown in red) and the model prediction/matching (shown in blue) components, plotting the total computation time as a function of the number of stairs in the staircase (ranging from 3 to 20). Both the prediction and matching times increase with the number of stairs, with the EKF component exhibiting a steeper rise. Fig. 8 illustrates the time required for stair surface segmentation, with times normalized for a point cloud of 250k points to account for varying point cloud densities across different staircases. Similar to the EKF, the computation time for segmentation shows an upward trend as the number of steps increases, reflecting the increased time required for a larger number of points. Despite these trends, the absolute segmentation times remain manageable, even for staircases with 20 steps. Crucially, the total time taken by both the estimation and segmentation methods, even for large staircases, remains under 30ms. This is well within 50ms budget to achieve our real-time execution target rate of 20Hz, confirming the feasibility of real-time operation.